Right now, Susan's age is 4 years more than four times Jane's age. In 20 years, Jane will be half of Susan's age then. How old are they now?
Thanks!

Answer by flame8855(424)

let s be the age of susan
let j be the age of jane
s=4+4j
j+20=(s+20)/2
solve
j= 8 years
s = 36 years
now jane is 8 years
and susane is 36 years